A copper wire having length L and diameter d has risistance R. What will the resistance be of a copper wire with length 2L and diameter 2d?


1
Expert's answer
2022-01-10T09:17:10-0500

Solution.

"R=\rho\dfrac{l}{S};"

"l_1=2l;"

"d_1=2d;"

"S=\pi \dfrac{d^2}{4};"

"S_1=\pi\dfrac{4d^2}{4}=\pi d^2=4S;"

"R_1=\rho\dfrac{2l}{4S}=\rho\dfrac{l}{2S}=\dfrac{R}{2}=\dfrac{1}{2}R;"


Answer:"\dfrac{1}{2}R."